OpenStudy (anonymous):

find the limit if it exists: lim x->3= (x-3)/(x^2-9)

OpenStudy (anonymous):

what i did was (x-3)/(x-3)(x+3)

OpenStudy (anonymous):

it became 1/(x+3) then i plugged in the 3 into the x and got 1/6, this is wrong though.

OpenStudy (helder_edwin):

no it is not. 1/6 is the right answer

OpenStudy (anonymous):

hang on

hartnn (hartnn):

1/6 is correct or verify the question.

OpenStudy (anonymous):

yea ur right who told u ur wrong

OpenStudy (anonymous):

oh wait i read the wrong answer in the back of my book. it is correct, thanks all!

OpenStudy (anonymous):

i wish i could give everyone a medal :(

OpenStudy (anonymous):

\[\lim_{x\to3}\frac{x-3}{x^2-9}=\lim_{x\to3}\frac{x-3}{(x+3)(x-3)}=\lim_{x\to3}\frac1{x+3}=1/6\]